What is the amount of money that Lebo has to invest, at a simple interest rate of 12%

 per year, to earn R5250,00

 interest after three years?


1
Expert's answer
2021-10-06T00:39:55-0400

"FV=PV(1+r)^n"


"FV-PV=PV((1+r)^n-1)=5250"


"PV=\frac{5250}{(1+r)^n-1}=\frac{5250}{1.12^3-1}=R12965.27"